OpenStudy (anonymous):

Find the distance between the two points. Round to the nearest tenth if necessary. (2, 5) and (–1, –5) A. 1 B. 109 C. 10.4 D. 3

OpenStudy (anonymous):

Use the distances formula: \[D = \sqrt{(x-x)^2 + (y-y)^2}\] So: 2 - -1 = 3, square that and you get 9 5 - -5 = 10, square that and you get 100. Add them together 100 + 9 = 109 and take the square root. D = 10.4403... So: C. 10.4